### v1

# A Hitchhikers Guide to Fine-Grained Face Forgery Detection Using Common Sense Reasoning

Abstract
Explainability in artificial intelligence is crucial for restoring trust, particularly in areas like face forgery detection, where viewers often struggle to distinguish between real and fabricated content. Vision and Large Language Models (VLLM) bridge computer vision and natural language, offering numerous applications driven by strong common-sense reasoning. Despite their success in various tasks, the potential of vision and language remains underexplored in face forgery detection, where they hold promise for enhancing explainability by leveraging the intrinsic reasoning capabilities of language to analyse fine-grained manipulation areas. As such, there is a need for a methodology that converts face forgery detection to a Visual Question Answering (VQA) task to systematically and fairly evaluate these capabilities. Previous efforts for unified benchmarks in deepfake detection have focused on the simpler binary task, overlooking evaluation protocols for fine-grained detection and text-generative models. We propose a multi-staged approach that diverges from the traditional binary decision paradigm to address this gap. In the first stage, we assess the models’ performance on the binary task and their sensitivity to given instructions using several prompts. In the second stage, we delve deeper into fine-grained detection by identifying areas of manipulation in a multiple-choice VQA setting. In the third stage, we convert the fine-grained detection to an open-ended question and compare several matching strategies for the multi-label classification task. Finally, we qualitatively evaluate the fine-grained responses of the VLLMs included in the benchmark. We apply our benchmark to several popular models, providing a detailed comparison of binary, multiple-choice, and open-ended VQA evaluation across seven datasets. https://nickyfot.github.io/hitchhickersguide.github.io/
1 Introduction
Recent developments in deep generative modelling have resulted in hyper-realistic synthetic images/videos with no clear visible artefacts, making the viewers question whether they can still trust their eyes. Unfortunately, despite its relevance in a wide range of applications, such technology poses a threat to society as it can be used for malicious activities [16]. In a world where synthetic images of a person, known as deepfakes, can easily be generated, it becomes crucial to fight misinformation not only by identifying manipulated images/videos in an automated manner but also by explaining the decision behind this classification to reinstate trust in Artificial Intelligence.
Numerous successful works on deepfake detection have been proposed in the literature to tackle the risks of face forgery [37, 76, 52, 7, 80]. Existing methods primarily rely on deep binary classifiers, resulting in black-box models that predict whether an input is real or fake. Consequently, explaining why those predictions are being made is not straightforward. To handle this issue, a few interpretable deepfake detection methods have been introduced by examining attention maps or weight activations [86, 71] to identify fine-grained areas of manipulation; however, these are based on a post-hoc analysis and thereby do not intrinsically incorporate an explainable mechanism. On the other hand, Vision Large Language Models (VLLMs) have emerged as a pioneering branch of generative Artificial Intelligence (AI), showcasing advancements in common sense reasoning and an inherent explainability that arises from the intrinsic nature of language [22]. They have demonstrated impressive capabilities in tasks such as Visual Question Answering (VQA) [47] and the generation of descriptive content for downstream applications [70], hence bridging the gap between vision-language understanding and contextual reasoning. However, despite these achievements, the explainable power of VLLMs remains under-explored in the field of deepfake detection, with only a handful of works exploring the vision and language capabilities for the binary classification of fake/real images [26, 8, 31, 69] all of which are evaluated on different benchmarks and metrics.
Explainable fine-grained detection –that is, identifying manipulation beyond the binary fake/real decision– in natural language is still in its infancy. However, as VLLM works for deepfake detection are expected to appear following the overwhelming success of foundation models in other tasks, two research questions need to be addressed: 1) “To what extent can existing VLLMs detect deepfake images and what rationale supports the decision?” and 2) “How do we fairly and comprehensively evaluate VLLMs in the fine-grained task?”. In deepfake detection, benchmarks have mainly focused on binary decisions and discriminative networks [81], making them unsuitable to answer these research questions. Indeed, they do not propose a unified method to match the generated responses to fine-grained categories. Similarly, existing benchmarks in Visual Question Answering (VQA) [19, 10] primarily address multi-class tasks, which may not be suitable for the multi-label nature of fine-grained deepfake detection as highlighted in [63].
In this work, our objective is to conduct a thorough quantitative and qualitative evaluation of VLLMs for the task of fine-grained deepfake detection in a systematic and scientific approach, employing a multi-stage protocol. In the first stage, we assess the models’ performance on the binary task using various prompts while also evaluating the model’s sensitivity to the provided instructions. In the second stage, we delve into fine-grained detection, aiming to identify areas of manipulation within a closed-form Visual Question Answering (VQA) framework, i.e. what areas from a given list are identified as manipulated. Subsequently, in the third stage, we extend our investigation by converting the fine-grained detection task into an open-ended question –that is, identifying areas without instructing the model to select from a list of categories. Here, as the task is a multi-label problem, we compare two matching strategies: a) using the cosine similarity between the generated text and ground truth labels and b) counting the occurrence of the class name in the generated text. Finally, we qualitatively evaluate the fine-grained responses generated by the VLLMs included in our benchmark, providing nuanced and new insights into their performance.
The main contributions of this work can be thus summarised as follows:
  • We introduce a novel approach in deepfake detection by proposing a method to transform the discriminative task into a Visual Question Answering (VQA) multi-label problem. To the best of our knowledge, this is the first work to do so in the field of deepfake detection, offering a fresh perspective on explainability through fine-grained analysis.
  • We establish a systematic, unified evaluation framework for current state-of-the-art (SOTA) VLLMs, facilitating consistent assessment across different models. This framework is designed to be extendable, ensuring fair and comprehensive comparisons with future models, thus promoting transparency and reproducibility in the evaluation process.
  • Through extensive comparison and analysis of the tested models and an ensemble of models, our study yields new insights into the capabilities and limitations of VLLMs in the context of deepfake detection. We will use these insights to advance research in the domain and hope to inform future developments and optimisations in model design and evaluation strategies.
2 Related Work
DeepFake generation encompasses various forms of facial manipulation, including face reenactment [6, 1], face swapping [24, 45], and entirely generated face images [32, 68]. Deepfake detection algorithms classify samples as real or fake [7, 80, 76, 37], relying on artefacts left by manipulation methods, often analysed qualitatively for explainability [21, 89]. However, this qualitative analysis happens on a secondary stage and primarily depends on human observers. While generative methods often use natural language instructions [54, 74, 58, 55, 77], explaining manipulations in natural language –a natural extension of the generation process to detection– is still an emerging area.
With the rise of VLLMs, recent works [8, 31, 69, 30, 72] explore vision and language for face forgery detection, primarily focusing on binary detection in a retrieval setting, with fewer [69, 52, 89] examining fine-grained areas usually as a secondary task. The latter have relied on generated pseudo-fake datasets to improve generalisation [69, 52, 89], which have a major drawback –that is, the use of pseudo-fake datasets hampers fair comparisons and does not reflect the current state-of-the-art in deepfake generation.
Several VLLMs foundation models [4, 15, 35, 44, 43], bridge the gap between vision and language. These are typically trained on very large datasets with general knowledge. As the computational and data resources needed to train VLLMs from scratch are very high, numerous works leverage the pre-trained networks in three main directions: a) exploring the latent feature space [14, 57] of vision and language, b) parameter efficient training in a downstream task [39, 73, 18] and c) evaluating foundation models in new domains [65, 75].
Benchmarks for classification tasks [19, 38, 2] in a VQA setting typically address the multi-class paradigm, which may not be appropriate for addressing explainability in DeepFake detection by adopting a multi-label fine-grained strategy, as several areas can be manipulated at once. A few preliminary works in DeepFake detection [3, 26] benchmark ChatGPT4111https://openai.com/gpt-4 and Gemini [^2] https://gemini.google.com; however, these have primarily focused on the more straightforward binary task and did not explore the reasoning capabilities of VLLMs for fine-grained labels. Furthermore, both these works focused on VLLMs that are not open-sourced with limited information regarding their training set and architecture; thus, it is not possible to assess whether the benchmarks are, in fact, zero-shot or whether they have been trained on deepfake-related image-language pairs. Within DeepFake detection, the vast majority of benchmarking works [60, 65, 46, 81, 34] have focused on binary discriminative networks and are therefore not fit to evaluate the capabilities of generative models such as VLLM, particularly for fine-grained labels.
In a nutshell, the main novelty of this benchmark compared to previous works [19, 81, 38, 2, 60, 65, 46, 34] is threefold: 1) it converts the multi-label classification task of face forgery detection to a VQA task so that VLLM’s common sense reasoning capabilities can be evaluated, 2) it systematically and consistently assesses VLLM capabilities on nine binary and three fine-grained benchmarks and 3) is offering an open source and extendable framework for future zero-shot or task-specific VLLMs, that ensures a fair comparison.
Figure 1: Overview of the proposed benchmarking method, using multiple stages to evaluate the performance of VLLMs in the context of deepfake detection. In the first stage (a), we assess the binary classification performance of VLLMs. In the second stage (b), we perform a fine-grained classification using multiple-choice instruction. In the third and final stage (c), we ask the model to identify fine-grained areas in open-ended VQA. The image example  [^3] Ground truth: Hair, Nose, Lip, Eyeis a sample from the SeqDeepFake dataset [63], and responses are generated using Llava-1.5 [43]
3 Common Sense Reasoning for Face Forgery Detection
Preliminaries: We formalise the language generation process of VLLM architectures, akin to standard VQA models, where the model is prompted with an image and a query to produce an auto-regressive answer. Given an image $\mathbf{X}_{v}\in\mathbb{R}^{H\times W\times C}$ and a text prompt $\mathbf{X}_{t}\in\mathbb{R}^{L\times d}$ as input, a sentence $\psi$ is generated represented as a sequence of word tokens. The generation can be represented by the function $p(\psi|\mathbf{X}_{v},\mathbf{X}_{t})=\prod_{j=0}^{|\psi|}p(\psi_{j}|\psi_{<j},\mathbf{X}_{v},\mathbf{X}_{t})$, where $H\times W\times C$ represent the image dimension, $L$ is the number of tokens, $d$ is the embedding dimension, ${\psi}=(\psi)_{0\leq j<|\psi|}$ is the generated sentence, and |.||.|| . | the cardinality. In VQA tasks, the model response aims to match human annotations. This task differs from typical classification problems due to the diverse semantic nature of questions and answers in natural language. The evaluation protocol is outlined for the binary case in Section 3.1 and for open-ended evaluation and multiple-choice of fine-grained labels in Section 3.2. An overview of the proposed method is shown in Figure 1.
3.1 Binary Classification to VQA
In binary classification, the task is to predict whether the image sample is a product of face forgery. We create a benchmark to assess VLLM capability in binary Deepfake detection by transforming the discriminative task into a VQA problem. We consider only the positive category for each image $\mathbf{X}_{v}$ to generate the relevant instruction; that is, we limit the prompt to identifying whether an image is a Deepfake and not whether it is a genuine sample. The prompt used is in the form:
[…]
where $\mathbf{s}_{i}\in\mathcal{S}$ denotes a set of standard terms used to describe deep fakes in the English language. The synonyms are employed to assess the reasoning capabilities of each tested model by investigating whether the understanding of the model is robust to the given instruction.
3.2 Fine-Grained Labels:
Fine-grained labels typically refer to manipulation areas. Predicting them necessitates the use of multi-label classification, as multiple areas can be manipulated at once. Following the initial binary prompt, a follow-up prompt to explain what areas of manipulation are identified is given to the VLLM with the same image, as shown in Fig. 1. We propose using two versions of follow-up prompts, one as an open-ended question and one as a multiple-choice. Specifically, the open-ended follow-up prompt follows the template:
[…]
For the multiple-choice instruction, we follow the template:
[…]
where $\mathbf{cls}_{i}\in\mathcal{C}$ is the class name of the $i$-th class from the set of target classes $\mathcal{C}$.
3.3 Matching Strategies:
To evaluate the generated responses, we employ several matching methods depending on the task. The stricter one uses an Exact Match (EM) approach, that estimates whether the generated sentence $\psi$ is exactly equal to the class name $cls_{i}$:
$p(\hat{y_{i}})=\begin{cases}1&\text{if $\psi\equiv\mathbf{cls}_{i}$}\\ 0&\text{if $\psi\neq\mathbf{cls}_{i}$}\end{cases}$ (4)
where $\hat{y_{i}}$ is the prediction for the $i$-th class. In the given task, an answer is considered correct only if the model output exactly matches the class names or ‘Yes’/‘No’ in the binary case. As the responses tend to be longer for fine-grained classification and reflect reasoning in natural language for a multi-label problem, a more appropriate matching strategy is to consider a response correct if the class name is Contained in the response, as proposed by Xu et al.  [75] – that is $p(\hat{y_{i}})=1,\text{if }\mathbf{cls}_{i}\in\psi$ and 00 otherwise. We extend this to include synonyms of class names, as several ways exist to describe some areas (e.g. “Bangs” could also be described as “Hair”). Finally, we propose adapting the text-to-text score (CLIP distance) proposed by Conti et al.  [14] for the multi-label task. This is done by using a sigmoid function over the cosine similarity matrix of the prediction embeddings and class name embeddings (obtained with a CLIP [59] text encoder), using an empirical temperature $t$ of $0.5$ so that $p(\hat{y_{i}})=\sigma(<\psi,\mathbf{cls}_{i}>\frac{1}{t})$. The symbol <.,.><.,.>< . , . > denotes the cosine similarity of the text embeddings and $\sigma(.)$ is the sigmoid function.
4 Experimental Set-Up
4.1 Tested VLLMs
We select four open-source state-of-the-art VLLMs to be included in this benchmark; specifically, we include LlaVa-1.5 [43] (an improved version of the LlaVa architecture [44]), BLIP2 [35] and finally InstructBlip [15] with Flan-T5 and Flan-T5-xxl language generators [13]. Finally, for the binary task, we include the CLIP [59] performance as a baseline and compare it against an ensemble of BLIP2 [35] and LlaVa-1.5 [43] following the ensembling strategies for VQA tasks [5], and GPT4v as an upper bound [^4] https://openai.com/index/gpt-4v-system-card/. Experiments using GPT4v are performed on a subset of 5k samples selected from each dataset and thus the results may be susceptible to some degree of bias from the sampling, which needs to be considered when comparing the models. The selection of the VLLM is guided by three factors. First, we select architectures with publicly available weights and training methods to ensure transparency and fairness in the evaluation. Second, we include methods that generate output in Natural Language rather than a set of features or classification predictions. Finally, we select methods that have achieved state-of-the-art performance on several zero-shot tasks. Additional model details, such as the number of parameters and pre-training information of the tested models, can be found in Appendix A.
4.2 Datasets
We evaluate the performance of our method on seven published challenging benchmarks and one pseudo-fake dataset; more specifically, seven datasets for binary detection and two for the fine-grained task. All are evaluated at the frame level, as in previous image works [79, 52, 49]. FF++: [60] consists of over 20k images of DeepFake images from 1000 videos, using four types of manipulation methods: Deepfakes, Face2Face, FaceSwap and NeuralTextures. The dataset is split into train, validation, and test with an 80%, 10%, and 10% split, respectively. DFDC: [17] is composed of 5k videos of real and manipulated faces split into 4,464 unique training clips and 780 unique test clips. Celeb-DF: [41] includes 590 original videos collected from YouTube with subjects of different ages, ethnic groups and genders, and 5639 corresponding manipulated videos. WildDeepFake: [88] is a challenging dataset for in-the-wild detection, which consists of 7,314 face sequences extracted from 707 videos that are collected completely from the internet. StyleGAN: Two sub-sets consist of curated images generated with StyleGAN3 [29] and StyleGAN2 [28] along with original ones for binary detection of facial images. SeqDeepFake: [63] dataset consists of 85k sequential manipulated face images based on two representative facial manipulation techniques, facial components manipulation [32] and facial attributes manipulation [27]. The labels include annotations of manipulation sequences with diverse sequence lengths. R-splicer: Augmenting real data by generating pseudo-fake images is a common practice in deepfake detection [36, 49, 11, 66, 87, 40]. Such methods simulate characteristic face-swap artefacts using simplistic operations on a predefined set of regions. In this work, we use a spliced dataset of 59k images to evaluate fine-grained labels of five regions –entire face, mouth, nose, eyes, eyebrows– as implemented by Mejri et al.  [49].
4.3 Metrics
Accuracy and the Area Under the Receiver Operating Characteristics (ROC) Curve (AUC) are the most common metrics used in DeepFake detection [81, 52, 65]. However, as the datasets in the task are massively imbalanced, we also use the harmonic mean of Precision and Recall (F1-score) for the binary task. Furthermore, we note that as AUC is a measure of the classifier’s performance at different thresholds, it has very limited value in the VQA task where matching strategies result in polarised decisions; however, we include it for reference. In the fine-grained task, we use mean Average Precision (mAP), AUC and F1-score as indicators of classification performance.
5 Results
(a) SeqDeepFake [63] attributes
(b) SeqDeepFake [63] components
(c) R-splicer dataset
Accuracy
(d) SeqDeepFake [63] attributes
(e) SeqDeepFake [63] components
(f) R-splicer dataset
AUC
(g) SeqDeepFake [63] attributes
(h) SeqDeepFake [63] components
(i) R-splicer dataset
F1-score
Figure 2: Exact Match (EM) Performance of each VLLM in terms of Accuracy (top), AUC (mid) and F1-score (bottom) for the top 3 synonyms “manipulated”, “synthetic” and “altered”
5.1 Binary Classification
Robustness to different prompts: We use seven synonyms for the positive class: “manipulated”, “deepfake”, “synthetic”, “altered”, “fabricated”, “face forgery” and “falsified”. As the binary task is simple and the instruction format is a ‘Yes’ or ‘No’ question, we use EM as a matching approach in this evaluation. In  Fig. 2, we see the performance of each tested model under the binary detection setting on the two sub-sets of SeqDeepFake [63] and the R-splicer dataset using the three best performing synonyms: “manipulated”, “synthetic” and “altered”. The first observation is that no VLLM clearly outperforms others across all datasets and metrics. However, we see that BLIP-2 [35] has the most robust performance to the given instruction, even though it is the smallest in terms of parameters. Furthermore, the additional parameters of T5-xxl [13] do not seem to aid the task compared to the base InstructBLIP [15] with T5 generator, as the base model performs comparably better across most benchmarks. We theorise that as the VLLMs have not been explicitly trained on image-language pairs of manipulated images, a large number of parameters on the language generation leads to more hallucinations [25, 78] for this simple but abstract task. Compared to the CLIP [59], models appear to have competitive performance with the exception of InstructBLIP with T5xxl LLM. When both base models, i.e. BLIP-2 [35] and LlaVa [43], have relatively good performance, the ensemble shows marginal improvement, particularly in terms of Accuracy and F1; however, this is not consistent therefore we do not continue the investigation to fine-grained labels. The detailed performance of all models and synonyms across all datasets and additional analysis on CLIP [59] features can be found in the Appendix.
(a) Accuracy
(b) AUC
(c) F1-score
Figure 3: Exact Match (EM) Performance of each VLLM on all nine benchmarks
Overall model performance: We average the performance of the three best-performing synonyms on all nine benchmarks in  Fig. 3. No model clearly outperforms others across all metrics and datasets; however, we can observe competitive performance from BLIP-2 [35] on the binary task, even though it is the smallest model in terms of parameters. We also see that all models struggle with the more challenging in-the-wild datasets, such as CelebDF [41], which highlights the need for further development to achieve adequate generalisation.
Vision Encoder Finetuning: We finetune contrastively the vision encoder of LlaVa on FF++ using a sigmoid loss [82] over an ensemble of prompts for the real/fake categories, and evaluate as described in the previous section. Training details for the vision encoder can be found in Appendix C. The architecture with the fine-tuned vision encoder shows improved within dataset and cross-dataset performance as shown in Tab. 1. Even without detailed captions or updating the LLM weights, we see there are still gains from a task specific vision encoder, particularly in terms of F1-score with an average improvement of nearly $4\%$ within dataset and nearly $2\%$ cross-dataset (for SeqDeepFake, CelebDF and StyleGAN2).
Table 1: Binary performance of LlaVa-1.5 [43] with a fine-tuned Vision Encoder against the zero-shot baseline.
LlaVa Baseline LlaVa w. fine-tuned Vision Encoder
Acc. F1 Acc. F1
FF++ 64.54 73.10 64.57 (+0.03%) 76.83 (+3.73%)
SeqDeepfake Attr. 58.92 66.15 61.22 (+2.30%) 68.03 (+1.88%)
SeqDeepfake Comp. 84.62 90.57 84.24 (-0.37%) 90.20 (-0.38%)
R-Splicer 87.11 92.50 87.31 (+0.20%) 92.62 (+0.12%)
DFDC 54.02 58.24 53.86 (-0.16%) 58.65 (+0.41%)
CelebDF 35.67 41.81 37.60 (+1.93%) 43.53 (+1.72%)
DFW 53.35 61.56 53.45 (+0.10%) 61.90 (+0.34%)
StyleGAN2 33.67 38.62 35.20 (+1.53%) 39.72 (+1.10%)
StyleGAN3 39.80 45.66 39.30 (-0.50%) 45.74 (+0.08%)
Metrics: In terms of the selected metrics, following the initial intuition, there is limited information we can get from the standard Accuracy and AUC used in the binary task. Both are heavily skewed by the label distribution, which is typically imbalanced in deepfake datasets; however, the latter may also not be fit for VLLMs as AUC measures performance at different thresholds, which are not present with EM and contains matching strategies. As such, we argue that for the task at hand, the F1-score –and consequently robust to imbalance metrics– are more appropriate.
5.2 Finegrained Evaluation
For the fine-grained task, we evaluate the performance of the selected models in the open and closed vocabulary settings. The fine-grained labels are evaluated on samples where the ground truth is positive – i.e., on DeepFake samples.
Open-Ended VQA: We first evaluate the selected VLLMs under the open vocabulary VQA setting on the three fine-grained datasets. The results using contains and CLIP distance matching are reported in Tab. 2(a) and Tab. 2(b) respectively. An EM strategy is not possible in multi-label tasks, so no such evaluation is performed. No model clearly outperforms others across all metrics and datasets. In fact, we can observe that, in most cases, they have comparable performance. This holds true for both contains and CLIP distance metrics. In terms of matching strategy, using the CLIP distance consistently and greatly improves recall, as is evident by the improvement in the F1-score and explicitly shown in Appendix H. This matching approach slightly lowers the mAP and AUC scores compared to the contains metrics; however, using the cosine distance to match the open-ended responses to the class categories semantically may offer a more reliable output for the class of interest, as seen by the F1-score.
Table 2: Model performance on open-ended fine-grained detection using 2(a)) contains and 2(b) CLIP matching
BLIP-2 InstructBLIP InstructBLIP-xxl LlaVa-1.5
mAP AUC F1 mAP AUC F1 mAP AUC F1 mAP AUC F1
SeqDeepFake [63] attributes $61.8$ $51.0$ $20.4$ $61.3$ $50.4$ $18.3$ $\mathbf{63.1}$ $\mathbf{53.6}$ $37.5$ $61.7$ $51.1$ $\mathbf{40.0}$
SeqDeepFake [63] components $59.5$ $50.5$ $14.7$ $59.2$ $50.0$ $4.1$ $\mathbf{60.2}$ $\mathbf{51.8}$ $\mathbf{26.2}$ $59.0$ $49.6$ $17.1$
R-Splicer $55.8$ $55.6$ $31.3$ $52.3$ $53.2$ $23.5$ $53.8$ $54.0$ $31.1$ $\mathbf{58.7}$ $\mathbf{57.5}$ $\mathbf{41.6}$
(a) Assessment of model performance during open-ended evaluation with contains distance matching.
BLIP-2 InstructBLIP InstructBLIP-xxl LlaVa-1.5
mAP AUC F1 mAP AUC F1 mAP AUC F1 mAP AUC F1
SeqDeepFake [63] attributes $\mathbf{63.0}$ $\mathbf{53.6}$ $73.5$ $59.9$ $50.9$ $74.0$ $60.4$ $50.7$ $55.5$ $61.0$ $51.3$ $\mathbf{74.1}$
SeqDeepFake [63] components $58.8$ $52.7$ $71.0$ $55.5$ $49.0$ $71.7$ $\mathbf{59.9}$ $\mathbf{55.7}$ $59.8$ $56.1$ $49.6$ $\mathbf{71.7}$
R-Splicer $54.3$ $55.3$ $66.2$ $48.5$ $49.3$ $66.5$ $54.0$ $53.1$ $60.3$ $\mathbf{56.7}$ $\mathbf{57.4}$ $\mathbf{66.5}$
(b) Assessment of model performance during open-ended evaluation with CLIP distance matching.
(a) mAP
(b) AUC
(c) F1
Figure 4: Assessment of model performance in multiple-choice settings, in terms of 4(a)) mAP, 4(b)) AUC and 4(c)) F1 during multiple-choice evaluation with contains matching.
Multiple choice VQA: The performance of the VLLMs on the multiple-choice instruction is shown in Fig. 4. Even though the open-ended setting is theoretically more challenging, the performances of all tested models are comparable to each other and worse on the multiple-choice instruction for both mAP and AUC. Regarding the F1-score, however, LlaVa [43] consistently performs better than other models. Under the multiple-choice setting, we observe that the models tend to mention all label names, which raises the number of False Positives –a significant limitation of the multiple choice setting– or respond with “All of them” or “None of them”, which makes matching of any sort more challenging and is reflected even more in the lower F1 score. Appendix G includes detailed metrics for each category.
5.3 Qualitative Evaluation
As the BertScore [85] is shown to correlate with human evaluation, we first present the Bert-score precision, recall, and F1 scores achieved by each model for the fine-grained open-ended responses compared with ground truth references that have been formatted using the prompt: “The areas that are $[\mathbf{s}_{i}]$ are $[\mathbf{c}ls_{i}]$”. The results of this evaluation, along with the score of human annotators [19, 10] on a subset of the R-Splicer dataset, are shown in Tab. 3. As in previous sections, no model clearly outperforms others across all benchmarks; however, we see that Llava-1.5 [43] has the most competitive performance for most benchmarks, closely followed by InstructBLIP [15]. This is consistent with qualitative evaluations on VQA tasks [19, 44].
Table 3: Open-ended qualitative evaluation with human annotators in Tab. 3(a) and BertScore [85] in Tab. 3(b)3(d)
Model Human Eval. Score
BLIP-2 $0.35$
InstructBLIP $0.36$
InstructBLIP-xxl $0.33$
LlaVa-1.5 $\mathbf{0.38}$
(a) Average score of Human Evaluation (R-splicer)
Model Precision Recall F1
BLIP-2 $79.77$ $78.75$ $79.24$
InstructBLIP $\mathbf{86.53}$ $83.22$ $84.81$
InstructBLIP-xxl $80.73$ $81.78$ $81.25$
LlaVa-1.5 $84.86$ $\mathbf{85.31}$ $\mathbf{85.08}$
(b) SeqDeepFake [63] attributes
Model Precision Recall F1
BLIP-2 $79.87$ $79.72$ $79.61$
InstructBLIP $81.12$ $83.89$ $\mathbf{86.90}$
InstructBLIP-xxl $82.57$ $81.77$ $81.01$
LlaVa-1.5 $\mathbf{87.40}$ $\mathbf{86.37}$ $85.39$
(c) SeqDeepFake [63] components
Model Precision Recall F1
BLIP-2 $79.55$ $79.76$ $80.04$
InstructBLIP $83.47$ $85.34$ $\mathbf{87.39}$
InstructBLIP-xxl $82.53$ $81.87$ $81.23$
LlaVa-1.5 $\mathbf{85.94}$ $\mathbf{86.33}$ $86.74$
(d) R-splice
Overall performance: In the simpler binary setting, BLIP-2 is more robust to instruction than other models with more parameters; however, when it comes to fine-grained evaluation, larger models show an advantage in reasoning and identifying areas of manipulation in the open-ended and multiple-choice settings. It is, however, worth noting that no model clearly outperforms others across all metrics and datasets. All of the results presented are based on zero-shot evaluations, where models are tested without being specifically trained for deepfake detection. Despite this, the models are able to leverage a semantic mapping between language and visual input from their very vast pre-training, giving them an inherent concept of "real" versus "fake." This capability suggests that these models possess some degree of understanding when it comes to identifying deepfakes. However, this general understanding falls far behind that of task-specific models. When we fine-tune the vision encoder, there is a notable improvement in performance. The vision-language models can better capture details and nuances in the input data, which enhances their deepfake detection capabilities. Nevertheless, due to the scarcity of high-quality captions and large-scale vision-language datasets tailored to deepfake detection, the improvements remain limited and only in the binary task. Overall, addressing these limitations—by creating specialised datasets and foundation models—could lead to substantial advancements in this area.
Limitations and Future Work: As the models in this work are all evaluated under zero-shot settings, their performance is below that of purpose-build networks seen in previous works [81], particularly for more challenging in-the-wild datasets. This further highlights the need for task specific models and more fine-grained deepfake datasets, which is a key finding of the experiments conducted in this work. A significant limitation is the lack of detailed language descriptions in datasets, making qualitative evaluation harder. Additionally, current datasets lack fine-grained labels, restricting assessments of manipulations to pseudo-fakes and SeqDeepFake [63]. Furthermore, as both the pertaining and evaluation datasets are not unbiased, the performance of all VLLMs is susceptible to the bias of the datasets, which is not addressed in this or previous benchmarks [81]. Identifying these shortcomings is important for future works on the task, particularly as VLLMs gain traction.
6 Conclusion
In conclusion, our proposed benchmark has several contributions; first and foremost, we propose a method to transform deepfake detection into a VQA problem beyond binary classification to leverage common sense reasoning as an inherent explainability mechanism. We show how this can be achieved in both a multiple-choice and open-ended VQA –with the latter being the most important use-case for new and unknown face forgery methods. This approach is used to evaluate a multi-label problem that is not typical of classic VQA. By doing so, we can systematically and consistently evaluate the common sense reasoning capabilities of current and future VLLMs in fine-grained deepfake detection.
Our selection of metrics and matching strategies allows for a fair evaluation of the proposed task. In particular, we include metrics that are robust to imbalance in both the binary and multi-label fine-grained tasks. Even though VLLMs in a zero-shot evaluation do not outperform purpose-built methods, the generated responses include reasoning, therefore holding promise for significant contributions in explainable deepfake detection, confirming the initial motivation behind examining the use of such models for the task and understanding the current capabilities. Moreover, as this benchmark can be extended in terms of models and datasets, it allows for a systematic and fair comparison of new language generation methods for explainable deepfake detection.
Ethics statement: The authors of this paper acknowledge the crucial role of ethical considerations in AI research and development. Our dedication lies in upholding principles of fairness and impartiality. Recognising the societal implications of generative technology (including VLLMs), we commit to transparency by openly communicating our findings and advancements with the research community.
Acknowledgments and Disclosure of Funding
This work is supported by the Luxembourg National Research Fund, under the BRIDGES2021/IS/16353350/FaKeDeTeR, and by POST Luxembourg. Experiments were performed on the Luxembourg national supercomputer MeluXina. The authors gratefully acknowledge the LuxProvide teams for their expert support.
Appendix
Appendix A Model Zoo
LlaVa-1.5 [43] is an extension of the LlaVa [44] model. We use the variant with CLIP-ViT-L-14 as a vision encoder and Vicuna-7b [12] language model. BLIP-2 [35] uses a QFormer architecture to bridge frozen language and vision encoders. We use the variant with CLIP-ViT-L-14 as a vision encoder and the 2.7b OPT [84] language model. InstructBLIP [15] is a family of VLLMs that exploits the basic BLIP-2 [35] architecture and advances the task by giving the instruction to both the QFormer and the LLM. We use two variants of the architecture, with two different LLMs from the T5 family [13]; in the base architecture, we use the CLIP-ViT-L-14 as a vision encoder and the T5-xl LLM. InstructBLIPxxl uses the same vision encoder and the T5-xxl language model. A comparison of all architectures in terms of parameters and pre-training datasets can be seen in Tab. 4. It is worth noting, that none of the pre-training datasets are related to deepfakes, making the task more challenging. For the ensemble, we chose BLIP-2 [35] and LlaVa-15 [43] , based of two main factors: a) they show the most competitive performance on most datasets as seen in Fig. 3 and b) they have the least overlap in terms of pre-training data thus we intuitively expect them having complementary information. The ensembling method adopted in this work is using score fusion with majority voting; in the occasions where the models disagree, the mean is taken. GPT4v is used for comparison in the binary tasks, however as the training details of this model are unknown, it should only be treated as an upper bound.
Architecture FLOPS # Params Pre-training Data
BLIP-2 [35] $0.38T$ $3.74B$ COCO [42], Visual Genome [33], CC3M [64], CC12M [9], SBU [56], and $115M$ images from the LAION400M [61]
InstructBLIP [15] $0.33T$ $4.02B$ COCO [42], WebCapFilt [35], TextCaps [67], VQAv2 [20], OK-VQA [48], AOK-VQA [62], OCR-VQA [51], LLaVA-Instruct-150K [44]
InstructBLIP-xxl [15] $0.53T$ $12.31B$ COCO [42], WebCapFilt [35], TextCaps [67], VQAv2 [20], OK-VQA [48], AOK-VQA [62], OCR-VQA [51], LLaVA-Instruct-150K [44]
LlaVa-1.5 [43] $4.14T$ $7.06B$ LLaVA-Instruct-150K [44], VQAv2 [20], OK-VQA [48], OCR-VQA [51]
Table 4: Comparisons of model FLOPS, number of parameters and pre-training datasets for selected VLLMs
Implementation Details. All experiments were conducted using four NVIDIA A100 GPUs, with 40GB of memory. We use the PyTorch deep learning framework for all model evaluation tasks and weights published on HuggingFace [^5] https://huggingface.co/.
Appendix B Binary Classification Prompts
Table 5: Reported cross-dataset performance of purpose-built discriminative SoTA models.
CelebDF [41] DFW [88] DFDC [17]
AUC mAP AUC mAP AUC mAP
UCF [80] 82.40 - - - 80.50 -
X-Ray [37] 79.50 - - - 65.50 -
Xception [60] 61.18 66.93 65.29 55.37 69.90 91.98
REECE [7] 70.93 70.35 68.16 54.41 - -
CORE [53] 74.28 - - - 73.41 -
LAA-Net [52] 86.28 91.93 57.13 56.89 69.69 93.67
The term deepfake is a colloquial term for a wide range of manipulations using generative models, from altering one small area all the way to fully generated images and videos. As such, the class name itself has several synonyms that can describe it. To assess the model’s robustness to instruction, we first prompt an LLM –specifically ChatGPT3.5 to give us synonyms for a deepfake. This is done as an automation step to incorporate the general consensus into the method without the author’s bias, following previous works [50]. The following synonyms are tested for the positive class: “manipulated”, “deepfake”, “synthetic”, “altered”, “fabricated”, “face forgery” and “falsified”. We show the detailed performance of all models on all synonyms in Tab. 6. To provide context, we also provide the cross-dataset performance of several discriminative SOTA works in Tab. 5. The models show the most consistent performance on synonyms “manipulated”, “synthetic” and “altered”; therefore, we do all subsequent analyses on these prompts.
Table 6: Binary Performance of tested models
Seq. Deepfake Attr. Seq. Deepfake Comp. R-Splicer FF++ DFDC CelebDF DFW StyleGAN2 StyleGAN3
Synonym Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1
manipulated 95.36 50.49 97.62 94.44 50.46 97.14 95.32 50.59 97.60 65.42 49.76 79.00 49.47 49.95 66.13 52.20 49.87 68.34 50.40 49.99 66.93 52.60 50.25 68.69 50.00 49.29 66.33
altered 96.06 50.25 97.99 94.62 49.93 97.23 95.85 50.22 97.88 65.76 50.01 79.24 49.50 49.98 66.21 51.60 49.31 67.73 50.37 49.96 66.94 53.20 50.84 69.13 51.49 50.76 67.66
synthetic 95.48 51.75 97.68 93.95 52.09 96.87 90.45 54.42 94.95 65.27 51.63 78.15 49.46 49.93 65.95 40.80 39.20 56.21 50.51 50.17 65.05 54.60 52.37 69.53 53.73 53.05 68.37
deepfake 93.92 51.54 96.86 93.15 54.50 96.43 89.47 54.91 94.40 64.66 52.19 77.28 49.35 49.82 65.81 39.40 37.79 55.24 50.48 50.16 64.18 53.80 51.53 69.16 50.75 50.07 66.33
fabricated 95.00 51.80 97.43 94.02 50.56 96.92 91.70 52.53 95.65 64.79 50.58 78.04 49.41 49.89 66.03 46.60 44.64 62.66 50.85 50.49 65.80 53.80 51.51 69.24 51.49 50.80 67.01
face forgery 96.11 49.98 98.01 94.72 51.87 97.28 94.18 51.18 97.00 65.03 50.42 78.36 49.44 49.92 66.01 53.00 50.69 68.79 49.67 49.27 66.19 53.60 51.26 69.31 51.49 50.76 67.66
falsified 95.58 52.10 97.73 93.29 51.74 96.52 92.03 53.25 95.83 65.42 51.46 78.38 49.40 49.88 65.94 42.80 41.07 58.55 52.92 52.56 67.06 53.80 51.57 68.99 52.99 52.30 68.02
(a) BLIP-2 [35] performance on the nine benchmark datasets.
Seq. Deepfake Attr. Seq. Deepfake Comp. R-Splicer FF++ DFDC CelebDF DFW StyleGAN2 StyleGAN3
Synonym Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1
manipulated 43.85 57.87 59.50 84.34 69.39 91.25 80.95 76.04 89.17 53.87 60.01 54.01 55.11 54.90 41.78 19.40 19.95 9.84 58.73 58.68 61.10 17.40 18.05 5.49 26.12 26.25 19.51
altered 45.10 55.92 60.86 82.10 65.58 89.91 81.31 76.41 89.40 54.32 61.01 53.83 55.51 55.29 42.39 19.40 19.94 10.24 58.37 58.32 61.23 17.80 18.41 6.80 27.61 27.70 23.62
synthetic 74.33 61.96 84.97 88.39 70.21 93.67 88.14 62.50 93.61 59.24 54.51 69.17 53.49 53.79 64.10 44.20 42.93 56.61 51.54 51.23 64.65 29.60 28.99 38.25 38.06 37.86 45.75
deepfake 20.19 58.62 29.42 62.94 75.23 75.93 66.76 78.77 79.26 49.04 60.62 39.09 53.54 53.17 24.19 26.80 28.13 0.54 59.07 59.30 43.77 26.80 28.13 0.54 28.36 28.74 4.00
fabricated 74.23 58.01 84.96 96.08 63.74 97.97 91.92 57.64 95.76 62.85 50.97 75.76 52.18 52.42 61.59 55.80 53.82 69.26 51.35 51.03 65.02 28.60 27.87 38.77 32.84 32.53 44.44
face forgery 19.90 58.47 28.99 54.55 74.75 68.60 71.10 82.75 82.42 51.26 62.73 42.42 54.76 54.42 30.14 26.80 28.13 0.54 56.81 57.03 41.85 27.00 28.32 1.08 32.84 33.29 4.26
falsified 19.62 58.33 28.55 57.76 75.13 71.51 64.94 81.28 77.80 49.44 61.70 38.24 54.15 53.80 27.19 26.00 27.29 0.54 55.86 56.14 34.03 26.00 27.29 0.54 29.10 29.52 2.06
(b) InstructBLIP [15] performance on the nine benchmark datasets.
Seq. Deepfake Attr. Seq. Deepfake Comp. R-Splicer FF++ DFDC CelebDF DFW StyleGAN2 StyleGAN3
Synonym Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1
manipulated 16.44 56.68 23.57 58.88 77.04 72.42 58.49 76.85 72.63 48.53 60.88 36.70 53.88 53.54 28.75 24.80 26.05 0.00 54.09 54.37 32.36 24.80 26.05 0.00 25.37 25.74 1.96
altered 17.21 57.08 24.80 60.98 78.14 74.19 60.99 77.24 74.72 49.04 61.25 37.78 53.99 53.66 29.59 23.40 24.52 1.54 54.90 55.15 36.37 23.20 24.33 1.03 38.81 39.39 0.00
synthetic 17.40 57.18 25.11 56.64 77.17 70.42 59.76 75.42 73.74 49.74 61.62 39.46 53.79 53.44 27.57 24.00 25.21 0.00 56.85 57.06 43.52 24.20 25.40 0.52 28.36 28.74 4.00
deepfake 7.98 52.29 8.77 13.43 54.42 16.24 41.08 69.28 56.04 42.22 56.31 22.40 52.28 51.85 14.48 35.00 36.76 0.00 51.36 51.73 13.75 35.00 36.76 0.00 38.81 39.39 0.00
fabricated 21.06 59.07 30.72 64.34 79.91 76.92 64.97 77.75 77.90 51.71 62.81 43.75 54.40 54.08 31.55 23.80 24.96 1.04 56.00 56.13 48.37 23.60 24.77 0.52 25.37 25.74 1.96
face forgery 9.42 53.04 11.47 30.63 63.48 42.46 41.83 69.76 56.84 42.83 56.74 23.90 52.24 51.82 16.20 34.20 35.92 0.00 51.66 52.02 15.64 34.20 35.92 0.00 26.87 27.25 2.00
falsified 16.73 56.83 24.04 58.74 78.28 72.25 63.55 78.75 76.76 49.68 61.77 38.94 53.81 53.47 28.49 24.40 25.57 1.56 54.25 54.49 35.95 24.00 25.19 0.52 26.87 27.25 2.00
(c) InstructBLIP [15] with Flan-T5-xxl [13] language model performance on the nine benchmark datasets.
Seq. Deepfake Attr. Seq. Deepfake Comp. R-Splicer FF++ DFDC CelebDF DFW StyleGAN2 StyleGAN3
Synonym Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1
manipulated 73.05 57.79 84.14 94.34 53.56 97.07 94.05 62.20 96.90 67.52 55.20 79.18 54.51 54.80 64.63 38.80 37.47 52.78 52.92 52.58 66.54 37.40 36.13 51.17 47.01 46.46 61.62
altered 89.16 55.66 94.22 94.30 51.65 97.06 96.10 52.45 98.01 66.42 51.41 79.42 50.67 51.11 65.95 49.00 46.85 65.31 51.05 50.65 67.17 47.20 45.12 63.74 50.75 50.04 66.67
synthetic 14.54 54.96 20.09 65.21 79.44 77.59 71.17 75.60 82.58 59.69 65.70 60.69 56.87 56.66 44.14 19.20 19.86 7.34 56.08 56.17 50.97 16.40 17.19 0.95 21.64 21.86 8.70
deepfake 8.00 52.16 8.29 38.46 67.53 51.91 51.03 74.43 65.99 46.16 59.25 31.44 54.18 53.79 23.32 34.60 36.33 0.61 53.63 53.98 20.52 34.20 35.92 0.00 37.31 37.88 0.00
fabricated 12.72 54.31 16.93 58.11 75.38 71.73 66.03 78.94 78.70 51.93 63.19 43.73 54.71 54.38 30.14 25.00 26.24 0.53 53.82 54.07 33.94 25.00 26.24 0.53 29.85 30.19 9.62
face forgery 12.62 54.56 16.72 56.78 76.25 70.50 71.86 84.60 82.94 55.75 66.12 50.36 56.38 56.05 33.71 29.40 30.79 2.75 55.02 55.29 35.12 28.60 30.02 0.56 32.84 33.31 2.17
falsified 16.85 56.46 23.86 64.83 77.98 77.32 78.19 84.70 87.30 60.02 68.31 58.50 56.98 56.72 40.85 23.00 23.93 5.87 57.44 57.62 46.59 21.40 22.42 1.50 26.87 27.18 7.55
(d) LlaVa-15 [43] performance on the nine benchmark datasets.
Seq. Deepfake Attr. Seq. Deepfake Comp. R-Splicer FF++ DFDC CelebDF DFW StyleGAN2 StyleGAN3
Synonym Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1 Acc. AUC F1
manipulated 95.38 58.70 97.63 94.83 67.61 97.34 96.24 54.77 98.08 66.33 74.05 79.69 49.61 54.59 66.18 52.40 51.21 68.60 50.50 59.84 67.05 53.79 58.98 68.39 50.75 50.00 67.33
altered 95.73 60.44 97.82 94.83 47.48 97.34 96.40 52.40 98.17 66.18 83.08 79.63 49.57 60.67 66.24 52.40 50.00 68.77 50.42 50.00 67.04 55.36 72.93 69.86 50.75 50.00 67.33
synthetic 80.63 54.15 89.02 65.31 56.09 77.70 88.48 53.20 93.83 63.24 56.20 74.45 51.09 52.60 63.29 44.20 39.67 57.14 48.37 46.35 61.71 58.20 58.93 66.07 55.22 64.14 68.42
deepfake 64.90 51.64 78.13 37.76 53.74 51.26 88.17 52.14 93.66 63.12 54.90 75.02 49.72 50.65 64.95 43.40 37.76 57.06 49.84 49.01 63.39 53.15 52.78 61.87 50.75 50.42 64.52
fabricated 94.62 52.63 97.23 13.17 51.97 18.13 91.13 53.02 95.33 64.09 56.23 75.85 49.80 50.66 64.39 51.80 50.20 65.02 49.81 48.39 64.50 55.05 55.36 64.60 55.97 65.32 68.78
face forgery 75.96 49.16 86.26 12.60 51.96 17.14 94.23 54.09 97.02 65.67 58.53 77.33 51.12 56.03 65.74 55.20 69.48 69.81 48.77 38.75 65.08 44.79 42.94 54.55 48.51 44.79 63.10
falsified 94.42 52.28 97.13 63.64 55.34 76.45 93.91 55.39 96.85 66.58 60.22 78.59 49.72 51.56 65.69 44.40 36.23 59.36 52.09 55.80 66.32 50.63 49.29 61.88 52.99 59.33 67.36
(e) Ensemble performance on the nine benchmark datasets.
Table 7: CLIP baseline performance on the binary task of the nine selected benchmarks.
Dataset Accuracy AUC F1
SeqDeepFake attributes [63] $54.62$ $63.46$ $69.63$
SeqDeepFake components [63] $35.80$ $49.10$ $50.38$
R-splice $17.14$ $44.33$ $25.98$
FF++ [60] $34.85$ $46.50$ $17.38$
DFDC [17] $46.84$ $46.75$ $40.93$
CelebDF [41] $59.80$ $60.95$ $49.11$
DFW [88] $49.67$ $49.90$ $31.09$
StyleGan2 [28]. $69.20$ $69.92$ $65.16$
StyleGan3 [29] $69.40$ $69.63$ $64.35$
(a) SeqDeepFake [63] attributes
(b) SeqDeepFake [63] components
(c) R-Spliced
(d) FF++
(e) DFW [88]
(f) DFDC [17]
(g) CelebDF [41]
(h) StyleGAN2
(i) StyleGAN3
Figure 5: t-SNE Visualisation of CLIP [59] image embeddings on the test set of the selected datasets (perplexity=50)
Appendix C Vision Encoder Fine-tuning
We fine-tune the CLIP-L/14-336 Vision encoder using LoRA [23] adaptors. Specifically, we add $32$ adaptors to the queries, keys, values and out projection of the attention heads, with an alpha of $32$, a dropout rate of $0.2$ and 4bit quantization. As the available datasets lack sample level descriptions, we use the text embeddings of the synonyms for the positive category and the embeddings of “real”, “original”, “unaltered”, “authentic”, “legitimate”, “genuine”, “bona fide” for the negative. We apply a Sigmoid loss over the cosine similarity [82] of all relevant text embeddings in the batch and train the vision encoder for 50 epochs. The updated weights are then used in the LlaVa-1.5 architecture.
Appendix D CLIP Embeddings
As most models use CLIP [59] variants as backbone vision encoders, we assess the separability of samples in real and manipulated images in each dataset by visualising them in a two-dimensional plane using t-SNE. The resulting visualisations can be seen in Fig. 5. The samples appear more separable in some datasets. Interestingly, the StyleGAN datasets, where the images are fully generated, seem to have more distinguishable latent representations from real images. The separability is not necessarily reflected in the language generation as seen in Sec. 5.1; to further examine the root cause of this, we first calculate the average image embedding of each class so as to create a class prototype and retrieve the top-10 nearest language token embeddings, seen in Tab. 8. The first observation that can be made is that none of the tokens seems related to the task at hand, which would potentially inform prompt selection, so without the reasoning capabilities of the LLM, the token retrieval on its own is not very informative. Secondly, we see that a number of tokens are repeated across datasets and for both classes, which can be attributed to the much smaller sub-space of the deepfake task compared to the CLIP latent space; so, while the image embeddings are somewhat separable for several datasets retrieving the nearest language tokens shows that the subspace is not very much related to face forgery. From these two observations, we can better understand the zero-shot performance of the tested foundation models. Finally, we show the performance of CLIP on the binary classification task in Tab. 7. We use an ensemble of prompts using the Imagenet prompt templates [59]; for the positive class, we average the embeddings of the prompts for synonyms “manipulated”, “synthetic” and “altered”; for the negative, we use “real”, “original” and “unaltered”. Contrary to previous VQA works [14, 19] that use CLIP retrieval as an upper bound, we see this is not the case in the task. This can be attributed to the more abstract definition of the class in the deepfake detection task, as well as the pre-training dataset, which is more object-oriented.
Table 8: Top 10 closest tokens to the class prototypes. Unique to the class tokens are in bold.
Dataset Original DeepFake
SeqDeepFake attributes natives, labeling, liz, saving, rink, demon, pitch, creole, godis, wentz natives, liz, labeling, rink, wentz, ronda, godis, creole, %), melissa
SeqDeepFake components natives, labeling, qld, romo, liz, creole, dhoni, ., hijab, gaining natives, qld, liz, gaining, romo, %), cuomo, klo, minions, cowgirl
R-splice natives, wentz, anglo, liz, %), qld, anca, romo, ronda, ural anglo, %), wentz, liz, natives, klo, anca, qld, ural, weed
FF++ natives, qld, wentz, anglo, cuomo, anca, liz, labeling, romo, melissa natives, qld, wentz, anglo, liz, cuomo, anca, %), weed, romo
DFDC natives, wentz, anglo, liz, ronda, qld, %), anca, romo, )!! anglo, wentz, natives, liz, %), ronda, qld, anca, )!!, romo
CelebDF natives, liz, wentz, ronda, %), anglo, qld, romo, labeling, anca anglo, natives, ural, klo, creole, %), weed, wentz, romo, minions
DFW anglo, %), ronda, wentz, ural, natives, liz, klo, rene, ator anglo, %), natives, ural, wentz, ator, ronda, klo, anca, qld
StyleGan2 natives, liz, wentz, ronda, %), anglo, qld, romo, labeling, anca natives, labeling, liz, rink, ronda, romo, creole, melissa, wentz, saving
StyleGan3 natives, liz, wentz, ronda, romo, labeling, aa, anca, %), ural natives, labeling, liz, ronda, rink, wentz, romo, saving, creole, melissa
Appendix E Human Evaluation
(a) Annotator Briefing
(b) Annotation Form
Figure 6: Example of Briefing( 6(a)) and Annotation Form( 6(b)) shown to human evaluators.
The human evaluation is based upon previous studies in VQA evaluation [19, 10]. The humans are asked to rate model predictions on a scale from 1 to 5, where 1 is completely wrong and 5 is completely correct.
Evaluation dataset: We use a subset of the pseudo-fake R-Splicer dataset for the human evaluation study, as it has artefacts visible to the human eye and is thus easier for the annotators to assess the response quality.
(a) Ground Truth: Mouth, Faceswap
(b) Ground Truth: Eyebrows, Eyes
(c) Ground Truth: Nose, Mouth, Faceswap
Figure 7: Samples of generated responses in an open-ended setting.
Human evaluation: We select 100 samples from the dataset and generate responses using the open-ended prompt. Each sample is annotated by the three human annotators on a scale from 1 to 5, where 1 is completely wrong, and 5 is completely correct. Each annotator is shown 50 samples. To reduce the workload on human annotators, we assign 1 (completely wrong) to all responses that describe the image content but no areas of manipulation. The annotations are then standardised between 0 and 1. An example of the form shown to human evaluators can be seen in Fig. 6.
Annotator agreement: as in previous works [19, 10] we use Krippendorff’s alpha to assess the inter-annotator agreement and obtain .75.75.75.75, which is a strong agreement given the complexity of the multi-label task. We average the annotator scores to receive the final gold standard for each sample and then obtain the average human score for each model.
Appendix F Qualitative Samples
By examining a few samples in  Fig. 7, we can also see that all models tend to hallucinate or provide a general description of the image content when they fail to identify specific areas of manipulation. Furthermore, while BLIP-2 [35] tends to respond more concisely –which is identified by Liu et al.  [44] as well–, we note that larger models attempt to provide a justification for their response, which is ultimately the goal of investigating VLLMs for this task.
As the available datasets lack sample specific descriptions of manipulation areas, we include BertScore of the tested VLLMs on the MagicBrush [83] dataset to assess response quality on the neighbouring task of image editing detection, that is language driven in table Tab. 9.
Table 9: BertScore of selected VLLMs on MagicBrush dataset
Precision Recall F1
Blip 85.44 84.37 84.57
InstructBlip 84.56 85.09 84.82
InstructBlip-xxl 81.07 84.55 81.88
LlaVa 1.5 83.08 84.77 83.87
Appendix G Multiple-choice VQA
Further to the analysis in Sec. 5.2, we show the performance of all models on the multiple-choice VQA setting in Tab. 10. In the multiple-choice setting, all models have comparable mAP and AUC, however, LlaVa-1.5 [43] shows a clear advantage in terms of Recall and F1.
Table 10: Performance of tested VLLMs in multiple-choice setting
AUC F1 mAP Recall
Bangs 49.8 0.0 60.3 0.0
Eyeglasses 49.9 8.0 63.4 4.6
Beard 49.8 1.0 61.0 0.5
Smiling 50.0 0.0 58.3 0.0
Young 50.0 0.0 62.3 0.0
Total 49.9 4.5 61.1 2.5
(a) SeqDeepFake attributes [63]
AUC F1 mAP Recall
nose 50.0 0.0 55.4 0.0
eye 50.8 10.2 66.8 5.5
eyebrow 51.0 11.0 58.3 6.0
lip 50.0 0.0 67.6 0.0
hair 50.4 9.7 48.9 5.4
Total 50.4 10.3 59.4 5.7
(b) SeqDeepFake comp [63]
AUC F1 mAP Recall
nose 51.3 46.9 52.5 42.2
eyebrows 50.4 57.0 52.0 63.1
eyes 50.2 56.2 51.5 61.9
mouth 49.9 45.4 51.9 40.9
faceswap 44.6 9.2 50.2 5.5
Total 49.3 42.9 51.6 42.7
(c) R-Splicer
BLIP-2 [35]
AUC F1 mAP Recall
Bangs 50.2 1.0 60.4 0.5
Eyeglasses 58.4 51.3 68.0 41.6
Beard 50.0 0.3 61.0 0.2
Smiling 50.0 0.0 58.3 0.0
Young 50.4 29.0 62.6 21.4
Total 51.8 20.4 62.1 15.9
(d) SeqDeepFake attributes [63]
AUC F1 mAP Recall
nose 50.0 0.0 55.4 0.0
eye 52.2 11.7 67.7 6.5
eyebrow 50.6 3.3 58.2 1.7
lip 50.0 0.0 67.6 0.0
hair 51.1 24.7 49.3 23.6
Total 50.8 13.2 59.6 10.6
(e) SeqDeepFake comp [63]
AUC F1 mAP Recall
nose 56.5 38.5 56.2 30.4
eyebrows 52.5 58.0 53.1 66.5
eyes 52.6 16.5 53.1 10.1
mouth 52.9 13.0 54.3 7.1
faceswap 50.0 0.0 51.6 0.0
Total 52.9 31.5 53.7 28.5
(f) R-Splicer
InstructBLIP [15]
AUC F1 mAP Recall
Bangs 50.4 18.5 60.7 13.2
Eyeglasses 54.3 75.9 65.7 90.3
Beard 50.0 0.1 61.0 0.1
Smiling 50.0 0.0 58.3 0.0
Young 50.0 0.0 62.3 0.0
Total 50.9 31.5 61.6 34.5
(g) SeqDeepFake attributes [63]
AUC F1 mAP Recall
nose 50.0 0.0 55.4 0.0
eye 57.3 39.1 70.3 31.2
eyebrow 54.4 21.0 60.8 13.0
lip 50.0 0.0 67.6 0.0
hair 50.1 38.4 48.8 44.4
Total 52.4 32.8 60.6 29.5
(h) SeqDeepFake comp [63]
AUC F1 mAP Recall
nose 53.4 59.0 53.6 64.6
eyebrows 52.7 31.5 53.3 22.9
eyes 51.0 8.5 52.0 4.7
mouth 57.0 41.6 56.5 32.3
faceswap 50.0 1.2 51.6 0.6
Total 52.8 28.4 53.4 25.0
(i) R-Splicer
InstructBLIP [15] with T5-xxl [13] LLM
AUC F1 mAP Recall
Bangs 52.2 38.1 61.4 27.3
Eyeglasses 58.2 73.0 67.5 77.8
Beard 57.1 74.3 64.7 86.7
Smiling 54.7 36.4 61.1 25.2
Young 49.5 74.1 62.1 92.6
Total 54.3 59.2 63.3 61.9
(j) SeqDeepFake attributes [63]
AUC F1 mAP Recall
nose 50.0 3.2 55.4 1.7
eye 48.8 76.2 65.9 91.0
eyebrow 49.2 71.8 57.4 96.0
lip 49.5 3.8 67.5 2.0
hair 52.2 65.4 49.9 94.9
Total 49.9 44.1 59.2 57.1
(k) SeqDeepFake comp [63]
AUC F1 mAP Recall
nose 49.6 67.6 51.6 98.2
eyebrows 50.1 67.9 51.8 98.4
eyes 50.2 67.7 51.5 98.9
mouth 50.0 68.0 51.9 98.4
faceswap 50.4 68.2 51.8 99.8
Total 50.1 67.9 51.7 98.7
(l) R-Splicer
LlaVa-1.5 [43]
Appendix H Open-ended VQA
The detailed performance of each model on the open-ended fine-graned detection can be seen on Tabs. 11, 12, 13 and 14. The matching strategy has a significant impact on Recall, as mentioned in Sec. 5.2, thus also increasing the F1 score.
Table 11: BLIP-2 [35] performance with contains and CLIP matching in open-ended VQA
AUC F1 mAP Recall
Bangs 53.9 33.7 62.5 22.2
Eyeglasses 56.4 40.3 67.0 28.0
Beard 50.1 2.1 61.1 1.1
Smiling 49.9 1.0 58.3 0.5
Young 44.6 24.9 60.3 18.2
Total 51.0 20.4 61.8 14.0
(a) SeqDeepFake attributes [63] with contains matching
AUC F1 mAP Recall
Bangs 56.0 73.5 64.1 99.9
Eyeglasses 60.0 75.9 69.6 99.9
Beard 59.9 74.1 66.8 99.9
Smiling 50.2 72.0 56.4 99.9
Young 51.6 75.1 61.2 100.0
Total 55.5 74.1 63.6 99.9
(b) SeqDeepFake attributes [63] with CLIP matching
AUC F1 mAP Recall
nose 49.4 20.1 55.2 13.2
eye 50.1 10.5 66.6 6.0
eyebrow 49.9 2.2 57.7 1.1
lip 50.6 11.6 68.0 6.5
hair 52.6 29.3 50.3 19.8
Total 50.5 14.7 59.5 9.3
(c) SeqDeepFake comp. [63] with contains matching
AUC F1 mAP Recall
nose 51.6 69.0 54.8 100.0
eye 49.8 77.3 64.1 99.8
eyebrow 50.4 70.9 55.7 100.0
lip 50.0 78.2 66.2 100.0
hair 49.7 63.2 46.8 99.8
Total 50.3 71.7 57.5 99.9
(d) SeqDeepFake comp. [63] with CLIP matching
AUC F1 mAP Recall
nose 55.8 31.6 55.9 20.5
eyebrows 50.4 2.5 52.0 1.2
eyes 54.0 37.3 53.7 27.8
mouth 52.7 20.5 53.7 12.4
faceswap 66.0 64.4 62.5 60.3
Total 55.8 31.3 55.6 24.5
(e) R-splicer with contains matching
AUC F1 mAP Recall
nose 54.4 66.6 58.5 99.9
eyebrows 51.7 66.6 52.7 99.9
eyes 53.5 66.3 53.7 100.0
mouth 52.8 66.8 54.1 100.0
faceswap 55.1 66.4 57.1 100.0
Total 53.5 66.5 55.2 100.0
(f) R-splicer with CLIP matching
Table 12: InstructBLIP [15] performance with contains and CLIP matching in open-ended VQA
AUC F1 mAP Recall
Bangs 51.3 9.1 61.1 5.2
Eyeglasses 50.5 3.0 63.7 1.6
Beard 50.2 1.2 61.1 0.6
Smiling 50.0 1.2 58.3 0.6
Young 50.0 76.8 62.3 99.9
Total 50.4 18.3 61.3 21.6
(a) SeqDeepFake attributes [63] with contains matching
AUC F1 mAP Recall
Bangs 49.6 73.6 58.5 99.6
Eyeglasses 47.2 75.7 60.5 99.1
Beard 58.0 73.9 64.7 99.3
Smiling 49.2 72.0 56.0 99.5
Young 50.5 74.9 60.0 99.5
Total 50.9 74.0 59.9 99.4
(b) SeqDeepFake attributes [63] with CLIP matching
AUC F1 mAP Recall
nose 50.1 0.9 55.5 0.4
eye 50.0 0.1 66.4 0.1
eyebrow 50.0 0.0 57.7 0.0
lip 49.7 0.4 67.5 0.2
hair 50.4 18.8 49.0 16.4
Total 50.0 5.1 59.2 4.3
(c) SeqDeepFake comp. [63] with contains matching
AUC F1 mAP Recall
nose 49.2 69.0 52.0 99.9
eye 46.9 77.3 61.7 99.8
eyebrow 49.5 70.8 54.1 99.9
lip 49.5 78.2 63.6 99.9
hair 49.6 63.3 46.1 99.9
Total 48.9 71.7 55.5 99.9
(d) SeqDeepFake comp. [63] with CLIP matching
AUC F1 mAP Recall
nose 54.0 16.8 55.2 9.3
eyebrows 50.4 3.3 52.1 1.7
eyes 50.4 9.3 51.6 5.4
mouth 53.0 19.4 54.0 11.8
faceswap 53.5 68.7 53.4 96.4
Total 52.3 23.5 53.2 24.9
(e) R-splicer with contains matching
AUC F1 mAP Recall
nose 50.8 66.6 52.0 99.9
eyebrows 49.2 66.6 49.3 99.9
eyes 49.9 66.3 49.2 99.9
mouth 51.0 66.8 51.0 100.0
faceswap 41.3 66.4 45.3 100.0
Total 48.5 66.5 49.3 99.9
(f) R-splicer with CLIP matching
Table 13: InstructBLIP [15] with T5-xxl LLM [13] performance, with contains and CLIP matching in open-ended VQA
AUC F1 mAP Recall
Bangs 54.4 27.4 62.9 16.7
Eyeglasses 60.6 65.6 69.1 59.7
Beard 51.0 6.3 61.6 3.3
Smiling 51.9 11.5 59.5 6.2
Young 50.0 76.8 62.3 100.0
Total 53.6 37.5 63.1 37.2
(a) SeqDeepFake attributes [63] with contains matching
AUC F1 mAP Recall
Bangs 45.8 46.3 56.2 40.7
Eyeglasses 56.1 65.3 65.7 65.9
Beard 55.0 60.3 63.4 59.2
Smiling 46.3 55.0 54.7 56.0
Young 50.2 50.6 61.9 43.7
Total 50.7 55.5 60.4 53.1
(b) SeqDeepFake attributes [63] with CLIP matching
AUC F1 mAP Recall
nose 50.1 0.9 55.5 0.4
eye 50.0 0.1 66.4 0.1
eyebrow 50.0 0.0 57.7 0.0
lip 49.7 0.4 67.5 0.2
hair 50.4 18.8 49.0 16.4
Total 50.0 5.1 59.2 4.3
(c) SeqDeepFake comp. [63] with contains matching
AUC F1 mAP Recall
nose 54.8 56.2 56.7 56.6
eye 58.2 57.7 68.5 49.4
eyebrow 51.2 56.1 55.1 56.8
lip 60.1 71.9 70.7 76.4
hair 54.1 57.3 48.3 70.4
Total 55.6 59.8 59.9 61.9
(d) SeqDeepFake comp. [63] with CLIP matching
AUC F1 mAP Recall
nose 53.1 18.4 54.0 10.6
eyebrows 50.5 7.5 52.1 4.0
eyes 52.1 38.5 52.5 29.9
mouth 53.7 29.4 54.2 19.2
faceswap 59.5 61.9 57.1 63.7
Total 53.8 31.1 54.0 25.5
(e) R-splicer with contains matching
AUC F1 mAP Recall
nose 53.5 57.2 53.5 63.7
eyebrows 52.7 59.1 51.9 69.6
eyes 51.4 57.2 50.7 66.0
mouth 54.3 62.2 53.7 78.3
faceswap 58.0 65.7 55.5 90.1
Total 54.0 60.3 53.1 73.5
(f) R-splicer with CLIP matching
Table 14: LlaVa-1.5 [43] performance, with contains and CLIP matching in open-ended VQA
AUC F1 mAP Recall
Bangs 48.9 56.9 59.9 55.9
Eyeglasses 53.6 39.0 65.2 27.4
Beard 52.5 16.7 62.5 9.5
Smiling 51.1 12.1 59.0 7.1
Young 49.3 75.3 62.0 96.1
Total 51.1 40.0 61.7 39.2
(a) SeqDeepFake attributes [63] with contains matching
AUC F1 mAP Recall
Bangs 46.4 73.5 56.9 99.9
Eyeglasses 51.9 75.9 65.9 100.0
Beard 57.9 74.1 64.0 100.0
Smiling 52.7 72.0 59.2 100.0
Young 47.5 75.1 58.8 100.0
Total 51.3 74.1 61.0 100.0
(b) SeqDeepFake attributes [63] with CLIP matching
AUC F1 mAP Recall
nose 48.8 19.7 54.9 12.7
eye 49.7 3.2 66.4 1.6
eyebrow 49.9 7.1 57.7 3.8
lip 50.3 12.3 67.8 6.8
hair 49.3 43.2 48.4 43.9
Total 49.6 17.1 59.0 13.8
(c) SeqDeepFake comp. [63] with contains matching
AUC F1 mAP Recall
nose 51.9 68.9 53.6 99.9
eye 50.0 77.4 63.1 100.0
eyebrow 48.7 70.8 54.1 100.0
lip 50.8 78.2 65.3 100.0
hair 46.8 63.3 44.5 100.0
Total 49.6 71.7 56.1 100.0
(d) SeqDeepFake comp. [63] with CLIP matching
AUC F1 mAP Recall
nose 57.6 34.8 57.4 22.5
eyebrows 53.5 19.5 54.3 11.2
eyes 57.4 43.2 56.2 33.3
mouth 55.9 36.7 55.8 25.6
faceswap 68.8 73.8 63.6 84.6
Total 58.7 41.6 57.4 35.5
(e) R-splicer with contains matching
AUC F1 mAP Recall
nose 56.8 66.6 60.0 100.0
eyebrows 54.5 66.6 56.2 100.0
eyes 56.5 66.3 56.0 100.0
mouth 55.0 66.8 55.6 100.0
faceswap 60.7 66.4 59.3 100.0
Total 56.7 66.5 57.4 100.0
(f) R-splicer with CLIP matching

### v2

@@
 Despite their success in various tasks, the potential of vision and language remains underexplored in face forgery detection, where they hold promise for enhancing explainability by leveraging the intrinsic reasoning capabilities of language to analyse fine-grained manipulation areas.
-As such, there is a need for a methodology that converts face forgery detection to a Visual Question Answering (VQA) task to systematically and fairly evaluate these capabilities.
-Previous efforts for unified benchmarks in deepfake detection have focused on the simpler binary task, overlooking evaluation protocols for fine-grained detection and text-generative models.
-We propose a multi-staged approach that diverges from the traditional binary decision paradigm to address this gap.
+For that reason, few works have recently started to frame the problem of deepfake detection as a Visual Question Answering (VQA) task, nevertheless omitting the realistic and informative open-ended multi-label setting.
+With the rapid advances in the field of VLLM, an exponential rise of investigations in that direction is expected.
+As such, there is a need for a clear experimental methodology that converts face forgery detection to a Visual Question Answering (VQA) task to systematically and fairly evaluate different VLLM architectures.
+Previous evaluation studies in deepfake detection have mostly focused on the simpler binary task, overlooking evaluation protocols for multi-label fine-grained detection and text-generative models.
+We propose a multi-staged approach that diverges from the traditional binary evaluation protocol and conducts a comprehensive evaluation study to compare the capabilities of several VLLMs in this context.
 In the first stage, we assess the models’ performance on the binary task and their sensitivity to given instructions using several prompts.
@@
 They have demonstrated impressive capabilities in tasks such as Visual Question Answering (VQA) [47] and the generation of descriptive content for downstream applications [71], hence bridging the gap between vision-language understanding and contextual reasoning.
-However, despite these achievements, the explainable power of VLLMs remains under-explored in the field of deepfake detection, with only a handful of works exploring the vision and language capabilities for the binary classification of fake/real images [26, 8, 31, 69] all of which are evaluated on different benchmarks and metrics.
+However, despite these achievements, the explainable power of VLLMs remains under-explored in the field of deepfake detection, with only a handful of works mostly exploring the vision and language capabilities for the binary classification of fake/real images [26, 8, 31, 70] all of which are evaluated on different benchmarks and metrics.
+To the best of our knowledge, Zhang et al.
+[87] is the only work employing a VQA approach in deepfake detection by proposing to extend the FF++ dataset with captions in natural language generated by humans.
+However, this work targets only one manipulated region at once, while deepfakes can incorporate several stacked manipulations [63, 68].
+Moreover, the provided augmented FF++ dataset does not allow for cross-dataset evaluation in a VQA setting without an extensive annotation effort, making it difficult to investigate the generalisation aspect.
+In addition, the fine-grained evaluation in previous works is limited as the more challenging open-ended VQA task is not explored.
 Explainable fine-grained detection –that is, identifying manipulation beyond the binary fake/real decision– in natural language is still in its infancy.
 However, as VLLM works for deepfake detection are expected to appear following the overwhelming success of foundation models in other tasks, two research questions need to be addressed: 1) “To what extent can existing VLLMs detect deepfake images and what rationale supports the decision?” and 2) “How do we fairly and comprehensively evaluate VLLMs in the fine-grained task?”.
-In deepfake detection, benchmarks have mainly focused on binary decisions and discriminative networks [81], making them unsuitable to answer these research questions.
-Indeed, they do not propose a unified method to match the generated responses to fine-grained categories.
+In deepfake detection, benchmarks have mainly focused on binary or multi-class decisions and discriminative networks [82, 87], making them unsuitable to answer these research questions.
+Indeed, they do not propose a unified method to match the generated responses to fine-grained multi-label categories.
 Similarly, existing benchmarks in Visual Question Answering (VQA) [19, 10] primarily address multi-class tasks, which may not be suitable for the multi-label nature of fine-grained deepfake detection as highlighted in [63].
-In this work, our objective is to conduct a thorough quantitative and qualitative evaluation of VLLMs for the task of fine-grained deepfake detection in a systematic and scientific approach, employing a multi-stage protocol.
+In this work, our objective is to conduct a thorough quantitative and qualitative evaluation of VLLMs for the task of fine-grained multi-label deepfake detection in a systematic and scientific approach, employing a multi-stage protocol without costly human captioning efforts.
 In the first stage, we assess the models’ performance on the binary task using various prompts while also evaluating the model’s sensitivity to the provided instructions.
-In the second stage, we delve into fine-grained detection, aiming to identify areas of manipulation within a closed-form Visual Question Answering (VQA) framework, i.e.
+In the second stage, we delve into multi-label fine-grained detection, aiming to identify areas of manipulation within a multiple-choice Visual Question Answering (VQA) framework, i.e.
 what areas from a given list are identified as manipulated.
@@
 The main contributions of this work can be thus summarised as follows:
-  • We introduce a novel approach in deepfake detection by proposing a method to transform the discriminative task into a Visual Question Answering (VQA) multi-label problem.
-To the best of our knowledge, this is the first work to do so in the field of deepfake detection, offering a fresh perspective on explainability through fine-grained analysis.
-  • We establish a systematic, unified evaluation framework for current state-of-the-art (SOTA) VLLMs, facilitating consistent assessment across different models.
-This framework is designed to be extendable, ensuring fair and comprehensive comparisons with future models, thus promoting transparency and reproducibility in the evaluation process.
+  • We introduce a novel evaluation protocol for deepfake detection under the Visual Question Answering (VQA) multi-label setting and without the use of human annotations.
+This is different from  [87] which is based on a succession of yes/no questions for fine-grained areas, resulting in a binary classification setting.
+It also relies on a relatively small dataset that cannot be extended without costly annotation efforts.
+In addition, our multi-stage protocol allows for open-ended VQA evaluation, which is a more challenging task.
+To the best of our knowledge, this is the first work to do so in the field of deepfake detection, offering a fresh perspective on explainability through fine-grained multi-label analysis.
+  • We present a systematic, unified evaluation study of current state-of-the-art (SOTA) VLLMs, facilitating consistent assessment across different models.
+This framework is designed to be extendable to any existing or future deepfake dataset, ensuring fair and comprehensive comparisons with future models, thus promoting transparency and reproducibility in the evaluation process.
   • Through extensive comparison and analysis of the tested models and an ensemble of models, our study yields new insights into the capabilities and limitations of VLLMs in the context of deepfake detection.
@@
 Benchmarks for classification tasks [19, 38, 2] in a VQA setting typically address the multi-class paradigm, which may not be appropriate for addressing explainability in DeepFake detection by adopting a multi-label fine-grained strategy, as several areas can be manipulated at once.
-A few preliminary works in DeepFake detection [3, 26] benchmark ChatGPT4111https://openai.com/gpt-4 and Gemini [^2] https://gemini.google.com; however, these have primarily focused on the more straightforward binary task and did not explore the reasoning capabilities of VLLMs for fine-grained labels.
+A few preliminary works in DeepFake detection [3, 26, 87] benchmark ChatGPT4111https://openai.com/gpt-4 and Gemini [^2] https://gemini.google.com; however, these have primarily focused on the more straightforward binary task and did not explore the reasoning capabilities of VLLMs for fine-grained labels.
 Furthermore, both these works focused on VLLMs that are not open-sourced with limited information regarding their training set and architecture; thus, it is not possible to assess whether the benchmarks are, in fact, zero-shot or whether they have been trained on deepfake-related image-language pairs.
+Zhang et al.
+[87] propose extending FF++ annotations with captions in natural language using human annotators.
+However, this method is limited to binary decisions, while a given deepfake image can undergo several manipulations.
+Furthermore, it does not explore the open-ended VQA setting and does not offer a method for cross-dataset evaluation without the need for a costly annotation process.
 Within DeepFake detection, the vast majority of benchmarking works [60, 65, 46, 82, 34] have focused on binary discriminative networks and are therefore not fit to evaluate the capabilities of generative models such as VLLM, particularly for fine-grained labels.
@@
 Finally, for the binary task, we include the CLIP [59] performance as a baseline and compare it against an ensemble of BLIP2 [35] and LlaVa-1.5 [43] following the ensembling strategies for VQA tasks [5], and GPT4v as an upper bound [^4] https://openai.com/index/gpt-4v-system-card/.
-Experiments using GPT4v are performed on a subset of 5k samples selected from each dataset and thus the results may be susceptible to some degree of bias from the sampling, which needs to be considered when comparing the models.
+Experiments using GPT4v are performed on a subset of 5k samples selected from each dataset, and thus, the results may be susceptible to some degree of bias from the sampling, which needs to be considered when comparing the models.
 The selection of the VLLM is guided by three factors.
@@
 We also see that all models struggle with the more challenging in-the-wild datasets, such as CelebDF [41], which highlights the need for further development to achieve adequate generalisation.
+Performance of GPT4v should be treated as an upper bound as we cannot assess whether the model has been trained on samples of the selected datasets.
+We see that GPT4v vastly outperforms the selected VLLMs on three benchmarks and has comparable performance on the rest, with the exception of FF++.
 Vision Encoder Finetuning: We finetune contrastively the vision encoder of LlaVa on FF++ using a sigmoid loss [83] over an ensemble of prompts for the real/fake categories, and evaluate as described in the previous section.
@@
 All of the results presented are based on zero-shot evaluations, where models are tested without being specifically trained for deepfake detection.
-Despite this, the models are able to leverage a semantic mapping between language and visual input from their very vast pre-training, giving them an inherent concept of "real" versus "fake." This capability suggests that these models possess some degree of understanding when it comes to identifying deepfakes.
+Despite this, the models are able to leverage a semantic mapping between language and visual input from their very vast pre-training, giving them an inherent concept of “real” versus “fake”.
+This capability suggests that these models possess some degree of understanding when it comes to identifying deepfakes.
 However, this general understanding falls far behind that of task-specific models.
 When we fine-tune the vision encoder, there is a notable improvement in performance.
-The vision-language models can better capture details and nuances in the input data, which enhances their deepfake detection capabilities.
+Vision-language models can better capture details and nuances in input data, which enhances their deepfake detection capabilities.
 Nevertheless, due to the scarcity of high-quality captions and large-scale vision-language datasets tailored to deepfake detection, the improvements remain limited and only in the binary task.
@@
 Limitations and Future Work: As the models in this work are all evaluated under zero-shot settings, their performance is below that of purpose-build networks seen in previous works [82], particularly for more challenging in-the-wild datasets.
-This further highlights the need for task specific models and more fine-grained deepfake datasets, which is a key finding of the experiments conducted in this work.
+This further highlights the need for task-specific models and more fine-grained deepfake datasets, which is a key finding of the experiments conducted in this work.
 A significant limitation is the lack of detailed language descriptions in datasets, making qualitative evaluation harder.